OpenStudy (anonymous):

A right triangle has base (x - 7) units and height (2x - 10) units. Part A: What is the square of the length of the hypotenuse of the triangle? Show your work. (4 points) Part B: What is the area of the triangle? (3 points) Part C: Using the solution obtained in Part B, explain the closure property of multiplication of polynomials. (3 points)

OpenStudy (anonymous):

@TrojanPoem

OpenStudy (trojanpoem):

hypo = k k^2 = (x-7)^2 + (2x -10)^ 

OpenStudy (anonymous):

So I have to find the other side, How do I do that ? I wrote exactly that down on paper but couldn't figure how to get what AB is

OpenStudy (trojanpoem):

No clue. I got 2 variables in one equation impossible to be solved.

OpenStudy (anonymous):

Alright, can you help me find the area ?

OpenStudy (trojanpoem):

yep Area = 0.5 * base * height

OpenStudy (anonymous):

How exactly do I multiply x - 7 and 2x - 10

OpenStudy (trojanpoem):

(x-7)(2x-10) x times 2x, -10 then -7 times 2x , -10

OpenStudy (anonymous):

Uh I have no idea was x times 2x is

OpenStudy (anonymous):

-10 x -7 = 70

OpenStudy (trojanpoem):

2x * x = 2x^2